OpenStudy (anonymous):

line AB = 5 centimeters, and line BC = squarroot 200 centimeters, what is the length of line segment AC

OpenStudy (raden):

diagram ?

OpenStudy (anonymous):



OpenStudy (anonymous):

use Pythagorean theorem

OpenStudy (raden):

use the pythagorean theorem : AC^2 = AB^2 + BC^2 AC^2 = 5^2 + (sqrt(200))^2 AC^2 = 25 + 200 AC^2 = 225 so, AC = sqrt(225) = ...

OpenStudy (anonymous):

so what would the length of AC be?